Please CLEARLY explain how the answer for this problem is c.

Table 11-5 A small island off the coast of Cape Cod contains two restaurants and two retail stores. Tourists need to take a ferry boat to reach the island, but with a recent slowdown in the economy, tourists are less willing to pay for the boat ride to visit the island. The owners of the restaurants and stores on the island-Restaurants 1 and 2, and Stores A and B_think that if tourists could ride the ferry for free, they would be happy to visit the island, eat and shop The business owners are considering contributing to a pool of money that will be used to pay for roundtrip ferry service each day. The table represents their willingness to pay, that is, the maximum amount that each business owner is willing to contribute, per day, to pay for each ferry trip. Store A $200 150 100 50 Store B Restaurant 1 $300 200 100 50 Restaurant 2 First ferry trip Second ferry trip Third ferry trip Fourth ferry trip S400 350 300 250 $180 90 30 7. Refer to Table 11-5. Suppose the cost to run the ferry for each roundtrip is S750 per day and the 4 business owners have agreed to split the costs of the ferry trips equally. Which of the following statements is correct? a. The owner of Restaurant 1 would prefer to have 3 ferry trips. b. The owner of Restaurant 2 would prefer to have 2 ferry trips. c. The owner of Store A would prefer to have 1 ferry trip d. All of the above are correct.
Answer #1

Answer 7

The correct answer is (c) the owner of Store A would prefer 1 ferry trip

Cost of each fairy ride is $750 and they all will split them equally. Hence each member will have to pay 750/4 = $187.5.

Hence They will prefer that number of ride such that there willingness to pay for last number of ride is greater than 187.5 i.e. Marginal Benefit > 0.

For Restaurant 1

Willingness to pay for 1st ferry trip = $300

Hence Marginal Benefit for one ferry trip is 300 - 187.5 = $112.5.

Willingness to pay for two ferry trip = $200

Hence Marginal Benefit for second ferry trip is 200 - 187.5 = $12.5

Willingness to pay for 3rd ferry trip = $100

Hence Marginal benefit for third ferry trip is 100 - 187.5 = -$87.5

Hence Restaurant 1 will prefer 2 two ferry trips.

Hence first option is wrong.

For Restaurant 2

Willingness to pay for 1st ferry trip = $400

Hence Marginal Benefit for one ferry trip is 400 - 187.5 = $212.5

Willingness to pay for two ferry trip = $350

Hence Marginal Benefit for second ferry trip is 350 - 187.5 = $162.5

Willingness to pay for 3rd ferry trip = $300

Hence Marginal benefit for third ferry trip is 300 - 187.5 = $112.5

Willingness to pay for 4th ferry trip = $250

Hence Marginal benefit for 4th ferry trip is 250 - 187.5 = $62.5

Hence Restaurant 2 will prefer at least four ferry trips.

Hence Second option is also incorrect

For Store A

Willingness to pay for one ferry trip = $200

Hence Marginal Benefit for first ferry trip is 200 - 187.5 = $12.5

Willingness to pay for second ferry trip = $150

Hence Marginal Benefit for second ferry trip is 150 - 187.5 = -$37.5

Hence Store A will will prefer onle 1 ferry trip

Hence third option is correct answer.

Hence, the correct answer is (c) the owner of Store A would prefer 1 ferry trip